Love Canal General Hospital wants to purchase a new blood analyzing device today. Its local bank is willing to lend it the money to buy the analyzer at a 3 percent monthly rate. The loan payments will start at the end of the month and will be $1,600 per month for the next eighteen months. What is the purchase price of the device?
Answer #1

Here, we can find out cash price using reverse calculation. We know at the end of 18 month the closing balance will be zero. Installment is $1600. That means interest for the 18th month is included in that installment. Interest rate is 3% monthly.

Therefore, interest included in 18th month = (1600/1.03)*0.03 = $46.60

That means principal is 1600 - 46.60 = $1553.40

Now for 17th month closing balance will be $1553.40

therefore, amount for interest calculation = $1600 + $1553.40 = $3153.40

So this amount includes 3% interest, therefore interest = (3153.40 /1.03) * 0.03 = $91.85

Principal paid in 17th month = 1600 - 91.85 = $1508.15

Similarly we can find the rest.

So at the end you will get Purchase price = $22,005.62

Month Opening Balance Principal Interest Installment Closing Balance 1,553.40 1,553.40 1,600.00 3,061.55 1,508.15 1,553.40 4,

If you want excel formulas:

1 Month Opening Balance Principal Interest Installment Closing Balance 2 18 =+C2 1600 =+C3+F3 1600 =+B2 =+E2-D2 =+E3-D3 =+E4-